what are the solutions of the equation? xsquared= 11x-24
Answer by stanbon(75887)

solutions of the equation?
xsquared= 11x-24
--------------------------
x^2-11x+24 = 0
---
x^2-8x-3x+24 = 0
x(x-8)-3(x-8) = 0
(x-8)(x-3) = 0
x = 8 or x = 3
=======================
Cheers,
Stan H.